OpenStudy (anonymous):

Medal A system of equations is shown below. 8x + 5y = 9 3x + 2y = 4 Part A: Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Part B: Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

I know I have to put it in y=mx+b form

OpenStudy (anonymous):

so -8 from both sides 5y=-8x+9

OpenStudy (anonymous):

2y=-3x+4 like that

OpenStudy (anonymous):

y=(4-3x)/2...y=2-1.5x, 8x +5*(2-1.5x)=9 x=-2

OpenStudy (anonymous):

y=2-1.5x=2+3=5

OpenStudy (anonymous):

y=(9-8x)/5

OpenStudy (anonymous):

x=-2,y=5

OpenStudy (anonymous):

multiply 2 *1--> 16x+10y=18 multiply 5*2--->15x+10y=20

OpenStudy (opcode):

@dg2 has it.

OpenStudy (anonymous):

is it correct